# Scaling Laws For Dense Retrieval

### v1

@@ 3.2. Model Architechture
With the development of large-scale pre-trained language models, advanced dense retrieval models in recent years have followed the Transformer’s structure. While some studies have explored using decoder-only architectures to generate dense vector representations of texts, mainstream dense retrieval models still employ encoder-only models such as BERT. Formally, a pre-trained Transformer, augmented with a projection layer, serves as the text encoder:
where represents the text input, denotes the parameters of the Transformer encoder, and and are the parameters of the projection layer.
@@
In our study, we experimented with Transformer models of various model sizes. With limited annotated query-passage pairs, it is usually difficult to train a large dense retrieval model from scratch. As a result, most dense retrieval models are initialized with pre-trained language models and then perform fine-tuning on the annotated data. Therefore, to align with prevailing research practices, we focus our analysis on fine-tuning dense retrieval models constructed with pre-trained language models in different sizes.
Previous studies have shown that different pre-training tasks could significantly affect the performance of dense retrieval models(Gao and Callan, 2021a, b; Izacard et al., 2021; Liu and Shao, 2022; Ma et al., 2022). To minimize such influence, we selected a series of models with identical pre-training configurations and only differ in parameter sizes. Specifically, for experiments on the English corpus, we chose 24 BERT checkpoints from the original Google release, with model sizes ranging from 0.6 million (BERT-Tiny) to 87 million parameters (BERT-Base)111https://github.com/google-research/bert. The model sizes are defined by non-embedding parameters in this paper.. For experiments on Chinese retrieval benchmarks, we selected the ERNIE series, which were pre-trained on Chinese corpora using tasks similar to BERT. To each model, we attached a projection layer to standardize the output vector (e.g., query and document vectors) dimensionality to 768 for consistent comparisons.
@@ 3.5. Evaluation Protocol
We now discuss how we evaluate the retrieval performance. The most widely adopted retrieval paradigm is to rank passages in the corpus based on the relevance scores predicted by the retrieval model and retrieve the Top-K candidates to form a ranked list. The performance of the retrieval model is then assessed based on the ranked list using well-defined ranking metrics such as NDCG@K and MAP@K. However, such metrics are not continuous due to their discrete nature and reliance on a cutoff parameter, K. Because the ranking metrics of a ranked list would not change unless the sequence of the passages changes, these ranking metrics are not sensitive to the changes of model outputs in many cases. Also, with the cutoff in ranking metric, a positive passage only contributes to the metric when ranked within the top K results. If it falls beyond K, whether at K+1 or further, it has no impact on the metric score. The characteristics of these existing ranking metrics make them unsuitable for the investigation of scaling laws in dense retrieval. To solve these problems, we propose to construct a more fluid and continuous metric that better reflects the overall retrieval capability of the models under various settings. Inspired by the analysis of scaling laws in large language models, which utilize the perplexity of token generations as evaluation metrics, we propose to use the contrastive perplexity as our evaluation metric. Formally, for each query-passage pair in the test set, we randomly select (256 in this paper) unlabeled passages and define the contrastive perplexity as:
We plot the IR metrics with the proposed contrastive perplexity for different models in Figure 2. As we can see from the figure, the correlation between the contrastive perplexity and existing ranking metrics is strong and positive. The correlation between contrastive perplexity and recall@1000, a popular metric used in most dense retrieval benchmarks, is particularly close to a linear correlation. Besides, the structure of contrastive perplexity is closely related to the training loss of dense retrieval models in our experiments. Therefore, we believe that using contrastive perplexity is an effective measure to assess the overall retrieval ability of models in our study.
@@ 4. Scaling Laws For Dense Retrieval
In this section, we show the results of our experiments and summarize our initial investigation of the scaling laws for dense retrieval.
@@ 4.1. Model Size Scaling
We first fine-tune models of various sizes using the human-annotated training pairs, while keeping all other settings identical. We utilized the entire training sets to simulate conditions of infinite data, in order to minimize the influence of performance constraints due to limited training data. We continuously monitored the contrastive perplexity in the test set and reported the best result of each model. It is important to note that this approach was adopted to mitigate the influence of suboptimal early stopping, which could lead to models being underfitted or overfitted.
In alignment with neural scaling law analysis(Kaplan et al., 2020), we define model size by the number of non-embedding parameters. Figure 3 illustrates the evaluation metric, namely the contrastive perplexity on the test set, with respect to model sizes. As shown in the figure, the retrieval performance improves (indicated by a lower test contrastive perplexity) as the model size increases. On the left side of the diagram, red stars represent the official checkpoints of variously sized BERT models, while blue points denote other official variants released concurrently. These variants differ in aspects such as the number of attention heads or feed-forward dimensions. The right diagram, in contrast, only features red stars, as the different shape variants of ERNIE are not publicly available.
Based on observation, we propose to fit the scaling law of dense retrieval models in terms of model sizes as follows:
where represents the number of non-embedding parameters of the model, and denotes the model’s contrastive perplexity on the test set. Parameters , and are the constants estimated based on the observations of different models’ performance.
Note that we introduce a parameter . This means that, when we have a sufficiently large model that has been adequately trained (setting to infinity), the contrastive perplexity approaches rather than zero. This is reasonable because, under the current settings of training and testing process, even an ideal model cannot reduce the contrastive perplexity to zero. Specifically, an important reason for this phenomenon is related to the annotation data in our experimental datasets. The contrastive perplexity used in our experiments directly reflects the distances between the model’s predictions and the actual human-annotated labels in the datesets. On one hand, human annotations are far from complete in MSMARCO and T2Ranking, leading to issues with false negatives. On the other hand, most existing retrieval datasets employ binary annotations, while an ideal model should fit a continuous relevance distribution. Therefore, even the best dense retrieval models cannot fit the human annotations perfectly, which explains why we need a bias term in Eq. (8)
We employed the commonly used least squares method to fit the linear curve, and the coefficient of determination (R²) suggests a good fit. Based on these results, we observe that the contrastive perplexity follows a power-law scaling in relation to the size of non-embedding parameters. The parameters of the fitted curve are detailed in Table 1.
@@
Such discoveries offer new perspectives for future research experiments. For example, given this scaling law, once a training dataset is available, we can initially train smaller models, fit the corresponding scaling curves, and then extrapolate them to predict the performance of larger models. This approach could significantly reduce the cost and effort of conducting experiments directly on larger models. More importantly, experimenting with different training strategies on smaller models can quickly validate the effectiveness of new approaches.
@@ 4.2. Data Size Scaling
We then fixed the model size and varied the size of the training data, defined by the number of annotated query-passage pairs. To minimize potential underfit problem caused by small models, we specifically fine-tuned the largest models in our experiment, i.e., the BERT-Base model. Here we only present the experiment results up to using all available annotatation data.
The results are plotted in Figure 4. Similarly, we fit our experiment observations with a log-linear curve. The coefficient of determination (R²) indicates a good fit. Based on these results, we infer that the contrastive perplexity follows a power-law scaling relative to the number of annotated query-passage pairs, with specific parameters detailed in Table 2
where represents the number of annotated query-passage pairs, and denotes the contrastive perplexity. Parameters , and are the constants to be determined.
@@
Orthogonal to our previous findings, this conclusion offers an alternative perspective for future dense retrieval experiments and the data annotation process. For instance, when dealing with a completely new corpus and determining the requisite amount of annotations, the traditional approach relies on past experience without a clear understanding of the sufficiency of data annotation. Now, by taking advantage of the data size scaling law, a potential approach involves initiating with a minimal amount of annotations, training a model, and fitting the corresponding scaling curve. Accordingly, we can approximate the necessary size of data annotation based on the target performance of the dense retrieval model. This method is more economical since, as observed, the loss diminishes almost linearly on a logarithmic scale with increasing data size, indicating diminishing returns on performance improvements given more data annotations. Therefore, while more data invariably enhances performance, excessive annotation may not be cost-effective. Predetermining the expected goals and corresponding data size through our observed scaling laws would be a more efficient and economical strategy.
5. Annotation Quality
@@
So far, we have observed strong scaling phenomena of dense retrieval model performance with respect to model sizes and data sizes. Yet, in the IR scenario, another aspect that remained unexplored is the quality of data annotations, which can significantly impact the effectiveness of the data. A pertinent question arises: Does the scaling effect hold true for data of different quality? To investigate this, we conducted experiments using annotations of different quality. Due to constraints in time and resources, our experiments were exclusively conducted on the MSMARCO Passage Ranking dataset. Specifically, we employed query generation techniques to create three distinct types of annotations:
Inverse Cloze Task (ICT): The unsupervised ICT method extracts sentences from passages and pairs them with their context to form query-positive pairs. Given its reliance on topic and semantic similarity within a passage, this method often includes noisy data, positioning it as the lowest-quality annotation.
@@
Supervised Generation Models: We utilize the well-trained docT5query to produce multiple queries for each passage, thereby forming query-positive pairs. The model is trained by human annotations, rendering it a higher-quality annotation method.
Large Language Models (LLMs): We instruct LLMs to generate relevant queries for given passages. Leveraging the substantial language comprehension and generation capabilities of LLMs, this method is considered to be of the highest quality. Practically, we used the recently open-sourced ChatGLM3, noted for its impressive performance in various downstream language tasks.
@@
For ICT and ChatGLM3, we generate a query for each positive document annotated by humans in the original datasets. For docT5query, we randomly sampled 500,000 passages from the corpus for query generation, since it is originally trained on the human annotated passages. In this way, we can align the training passages with human annotations and other annotations. Also, it’s important to note that, despite employing different data generation methods, our evaluations consistently utilized the human-annotated development set. The results are reported in Figure 5.
@@
From the figure, we observe that the retrieval performance scales with respect to different annotation qualities. Comparing the three methods of query generation, it is evident that, on the log-linear curve, ICT exhibits the smallest slope. This observation aligns with our expectations. ICT, as a weak supervision method, involves considerable noise, thus limiting the enhancements in retrieval tasks when we increase the data size. Therefore, previous studies usually use such weak supervision only in the retrieval-oriented pre-training phase, not the fine-tuning phase. The data quality from docT5query is better, but not as good as human annotations, which could be attributed to the limited generation ability for a relatively small T5-base model. The data from ChatGLM3 is of the highest quality, which results in a more steep slope in Figure 5.
Additionally, we include human-annotated data for reference and notice an interesting phenomenon in Figure 5: the scaling slope of models trained with ChatGLM3 generated data is larger than the slope of human annotations. This observation indicates that, with larger annotation sizes, the performance of dense retrieval models trained purely based on the data generated by ChatGLM3 might potentially outperform the models trained with human annotations, which shows the potential of LLMs in terms of data augmentation in retrieval tasks. Yet, it’s crucial to note that it’s not entirely appropriate to directly compare the curve for human-annotated data with the other curves. On the one hand, besides docT5query, we select the human-annotated positive passages in MSMARCO as candidates when generating queries, which may benefit the query generation methods in terms of data collection because low-quality documents are naturally excluded from the annotation process. On the other hand, the nature of human annotation – labeling relevant passages for a given query – differs inherently from query generation, which involves generating a relevant query for a given passage. Still, the results in Figure 5 reveal the potential of LLM-based data annotation for dense retrieval, which is worth more investigations in future studies.
@@ 6. Application in Budget Allocation
In this section, we showcase a potential application of the scaling laws for dense retrieval observed in our experiments. To get started, we first describe how to combine the above observations into a single function between model performance and model/data sizes. Inspired by the scaling laws of LLMs (Kaplan et al., 2020), on MSMARCO, we employ the following equation to describe the scaling effect as:
Similarly, represents the model size and data size, respectively. are the fitting parameters. We employed results with different model sizes and data sizes to fit the parameters, which are used to predict the remaining data points. Figure 6 illustrates the contrastive perplexity across different resource settings. In this figure, solid dots represent the data used for curve fitting, while the dashed line indicates the resulting fitted curve. The red stars denote data points utilized to evaluate the accuracy of our predictions.
We subsequently attempt to estimate the comprehensive cost associated with the lifecycle of dense retrieval models, including data annotation, model training, and model inference. The total cost of training a model with parameters using data points is given by:
@@
To better illustrate the relationship between the predicted contrastive perplexity and the total cost, we provide approximate values for the cost factors as follows. The cost of human annotations is approximated at $0.6 per query-passage pair(Althammer et al., 2023). For computational costs, according to previous studies(Kaplan et al., 2020; Clark et al., 2020), the training and inference computation for Transformer can be assumed by and FLOPs, respectively. We refer to common cloud computing and the price for using an A100 80G GPU is assumed to be $3.93 per hour222From https://cloud.google.com/compute/gpus-pricing, the cost for A100 80G GPU is $2867.5 per month., with the peak computational power around 312 TFLOPs. For the training phase, we assume that the model is trained for 10,000 steps on a single A100 GPU. At each step, the model encodes a query, a positive passage and a negative passage with a batch size of 256. Each query is around 30 tokens and each passage is around 60 tokens. For the inference phase, we assume that the model is employed in a web search engine. Based on public statistics, we assume that there are around 30 trillion web pages in Google’s index333From https://en.wikipedia.org/wiki/Google_Search, the estimated size of Google’s index is around 30 trillion in 2012.. The inference cost for a dense retrieval model predominantly involves encoding the entire corpus. We estimate that each web page contains approximately 512 tokens. We assume the GPU utilization efficiency is 25%, then we have
@@
To get started, our first analysis excludes the cost of inference and focuses on annotation and training. Figure 7 shows the predicted contrastive perplexity against model size under different cost budget. It is clear that for a fixed cost budget, as the training model size increases, the predicted retrieval performance initially improves and then deteriorates. With relatively small models, despite that we have a sufficient budget for data annotation, the capacity of the dense retrieval model would limit its overall performance. Conversely, when training models with large parameter sizes, the budget constraints for data annotation limit the amount of data that can be used for training, leading to suboptimal retrieval performance. Also, Figure 7 indicates that, without the consideration of inference cost, the optimal model size grows with an increased budget and can be over 13 billion parameters, which is not always practical in real-world scenarios. This is primarily due to that, compared to the cost of model training, human annotation is significantly more expensive. Therefore, under a limited budget, maximizing model size can yield better results.
Now, after including the inference costs, we have the analysis results in Figure 8. It is clear that the optimal model size significantly diminishes, resulting in models with million-scale parameters, even under a larger budget. This observation suggests a notable distinction between dense retrieval and language generation tasks: the corpus size for retrieval is often large, making the encoding cost prohibitively high when opting for billion-scale models.
7. Conclusions and Future Work
In this paper, we have focused on investigating the scaling laws for dense retrieval models. We conducted extensive experiments with varying model sizes, data sizes, and annotation data. Our key finding is that in dense retrieval tasks, when employing the proposed contrastive perplexity as the evaluation metric, retrieval performance exhibits a precise power-law scaling in relation to both model size and data size. Moreover, we further investigate the influence of different data augmentation methods and showcase potential applications of the fitted scaling law. However, there are limitations to this work. Firstly, due to limited time and computation resources, our experiments only involve BERT-series models, and the variations in model size and data size are conducted within a relatively small range. Therefore, whether the corresponding scaling laws hold true for models on the scale of billion-scale models requires further investigation. Secondly, different training strategies, training steps, and the corpus size of datasets are factors that could affect scaling. These aspects were not explored in this study and are left for future work.

### v2

@@ 3.2. Model Architechture
With the development of large-scale pre-trained language models, advanced dense retrieval models in recent years have followed the Transformer’s structure. While some studies have explored using decoder-only architectures to generate dense vector representations of texts, mainstream dense retrieval models still employ encoder-only models such as BERT due to its bi-directional modeling ability. Formally, a pre-trained Transformer, augmented with a projection layer, serves as the text encoder:
where represents the text input, and and are the parameters of the projection layer.
@@
In our study, we experimented with Transformer models of various model sizes. With limited annotated query-passage pairs, it is usually difficult to train a large dense retrieval model from scratch. As a result, most dense retrieval models are initialized with pre-trained language models and then perform fine-tuning on the annotated data. Therefore, to align with prevailing research practices, we focus our analysis on dense retrieval models initialized from different sizes of pre-trained language models.
Previous studies have shown that different pre-training tasks significantly affect the performance of dense retrieval models (Gao and Callan, 2021, 2022; Izacard et al., 2022; Xiao et al., 2022; Ma et al., 2022). To minimize such influence, we select a series of models with identical pre-training configurations and only differ in parameter sizes. Specifically, for experiments on the English corpus, we chose 24 BERT checkpoints from the original Google release (Devlin et al., 2019), with model sizes ranging from 0.5 million (BERT-Tiny) to 82 million parameters (BERT-Base)222https://github.com/google-research/bert. Following Kaplan et al. (2020), we define the model size as the number of non-embedding parameters.. For experiments on Chinese retrieval benchmarks, we selected the ERNIE series (Sun et al., 2019), which were pre-trained on Chinese corpora using tasks similar to BERT. To each model, we attach a projection layer, as shown in Eq. (2), to map the output dimensionality of embeddings to 768 for consistent comparisons.
@@ 3.5. Evaluation Protocol
We now discuss how we evaluate the retrieval performance. The most widely adopted retrieval paradigm is to rank passages in the corpus based on the relevance scores predicted by the retrieval model and retrieve the Top-K candidates to form a ranked list. The performance of the retrieval model is then assessed based on the ranked list using well-defined ranking metrics such as NDCG@K and MAP@K. However, such metrics are not continuous due to their discrete nature and reliance on a cutoff parameter, K. Because the ranking metrics of a ranked list would not change unless the sequence of the passages changes, these ranking metrics are not sensitive to the changes of model outputs in many cases. Also, with the cutoff in ranking metric, a positive passage only contributes to the metric when ranked within the top K results. If it falls beyond K, whether at K+1 or further, it has no impact on the metric score. The characteristics of these existing ranking metrics make them unsuitable for the investigation of scaling laws in dense retrieval.
To solve these problems, we propose to utilize a continuous metric that sensitively reflects the overall retrieval capability of the models. Inspired by the analysis of scaling laws in large language models, which utilize the perplexity of token generations as evaluation metrics, we propose to use the contrastive entropy as our evaluation metric. Formally, for each query-passage pair in the test set, we randomly select a fixed number (256 in this paper) of negative passages and define the contrastive entropy as:
We investiagte the correlation between the contrastive entropy and existing ranking metrics. We train multiple dense retrieval models. To efficiently evaluate their retrieval performance, we sample a subset corpus that contains 100,000 passages during evaluation. Figure 2 shows the contrastive entropy and ranking metrics, including MAP@10, NDCG@10, and Recall@1000. We can see that the correlation between the contrastive entropy and existing ranking metrics is strong and positive. It is close to a linear correlation. Therefore, we believe that using contrastive entropy is an effective measure to assess the overall retrieval ability of models in our study.
Figure 2 also shows a critical point around contrastive entropy, where the top ranking performance evaluated with traditional metrics substantially improves. We attribute this phenomenon to emergent ranking ability. Concurrently, Du et al. (2024) also observe this phenomenon in generation tasks. They find emergent abilities are tightly related to a certain loss value. We leave further exploration to future studies.
@@ 4. Scaling Laws For Dense Retrieval
In this section, we show the results of our experiments and summarize our initial investigation of the scaling laws for dense retrieval. Specifically, we aim to thoroughly investigate the following three research questions:
How does model size impact dense retrieval performance?
How does annotated training data size influence dense retrieval performance?
@@ 4.1. Model Size Scaling
We finetune models of various sizes using the human-annotated training pairs. The finetuning is performed on the entire training sets. We do not utilize early stopping and instead report the best test set loss throughout the training process. This is mainly to mitigate the influence of suboptimal early stopping, which could lead to models being underfitted or overfitted.
Figure 3 illustrates the contrastive entropy on the test set with respect to model sizes. As shown in the figure, the retrieval performance improves (indicated by a lower test loss) as the model size increases. On the left side of the diagram, red stars represent the official checkpoints of variously sized BERT models, while blue points denote other official variants released concurrently. These variants differ in aspects such as the number of attention heads or feed-forward dimensions. The right diagram, in contrast, only features red stars, as the different shape variants of ERNIE are not publicly available.
Based on the observation, we propose to fit the scaling law in terms of model sizes as follows:
where represents the number of non-embedding parameters of the model, and denotes the model’s contrastive entropy on the test set. Parameters , and are the coefficients.
Note that we introduce a parameter , which represents a irreducible loss term. It means that a sufficiently large model (setting to infinity) can only reduces the loss to rather than zero. This irreducible loss is reasonable given the incomplete annotations and subjective understanding of relevance. On one hand, some relevant passages may not be annotated because they are not sucessfully recalled and are outside the annotation pool. On the other hand, relevance may be subjective to different annotators, which results in even imperfect agreement among different human annotators. Consequently, it is hard for models to perfectly agree with human annotations. Therefore, we believe there should be a irreducible term in the scaling law.
We employ least squares method to fit the linear curve. The coefficients are detailed in Table 1. The coefficient of determination (R²) suggests a good fit. Based on these results, we validate that the contrastive entropy follows a power-law scaling in relation to the size of non-embedding parameters.
@@
Such discoveries offer new perspectives for future research experiments. For example, given this scaling law, we can initially train smaller models, fit the corresponding scaling curves, and then extrapolate them to predict the performance of larger models. This significantly reduces the cost of conducting experiments directly on larger models and instead offers the opportunity to experiment with different training strategies on smaller models to validate the effectiveness of new approaches.
@@ 4.2. Data Size Scaling
We then fix the model size and vary the size of the training data, defined by the number of annotated query-passage pairs. To minimize potential underfit problems caused by small models, we finetune the largest model in this experiment, i.e., the BERT-Base model. Here we present the experiment results up to using all available annotation data.
The results are shown in Figure 4. Similarly, we fit the scaling law in terms of data size with the following log-linear curve:
where represents the number of annotated query-passage pairs, and denotes the contrastive entropy. , and are coefficient to be estimated. The coefficient of determination (R²) indicates a good fit. Based on these results, we infer that the contrastive entropy follows a power-law scaling relative to the number of annotated query-passage pairs, with specific parameters detailed in Table 2.
@@
This finding offers an alternative perspective for future annotation process. For instance, to determine the amount of annotations for a new corpus, the traditional approach relies on past experience without a clear understanding of the sufficiency of data annotation. With the data-size scaling law, a potential approach is initiating with a minimal amount of annotations, training a model, and fitting the corresponding scaling curve. Accordingly, we can approximate the necessary size of data annotation based on the target performance of the dense retrieval model. This approach establishes a clear relationship between data annotation and the desired performance outcomes. It allows researchers to have a precise expectation of future model performance, facilitating more effective planning and budgeting for annotation tasks.
4.3. Annotation Quality
@@
So far, we have observed strong scaling phenomena of dense retrieval model performance with respect to model sizes and data sizes. Yet, in the IR scenario, another aspect that remained unexplored is the quality of data annotations: Does the scaling effect hold true for data of different quality?
To investigate this, we conduct experiments using annotations of different quality. Due to constraints in time and resources, our experiments are exclusively conducted on the MSMARCO Passage Ranking dataset. We employ query generation techniques to create three distinct types of annotations:
Inverse Cloze Task (ICT): ICT extracts sentences from passages and and uses the sentence as pseudo-query for the passage. Since it ignores the semantic information, the generated data is of low quality.
@@
Supervised Generation Models: We utilize docT5query (Nogueira et al., 2019a) to produce multiple queries for each passage. DocT5query is trained on human annotations. The generated data is of higher-quality than ICT’s.
Large Language Models (LLMs): We instruct LLMs to generate relevant queries for given passages. Since LLMs are strong in language understanding and generation, we consider the data quality to be better than both ICT and docT5query. We adopt ChatGLM3 (Zeng et al., 2023) due to its impressive performance in various tasks. The prompt for query generation is shown in Appendix A.
@@
For ICT and ChatGLM3, we generate a query for each positive document annotated by humans in the original datasets. For docT5query, we randomly sampled 500,000 passages from the corpus for query generation, since it is originally trained on the human annotated passages. In this way, we can align the training passages with human annotations and other annotations. Also, it’s important to note that, despite employing different data generation methods, our evaluations consistently utilize the human-annotated development set. The results are reported in Figure 5.
@@
We can see that the retrieval performance scales with respect to different annotation qualities. Comparing the three methods of query generation, the log-linear curve of ICT exhibits the smallest slope. This observation aligns with our expectation that ICT is a weak supervision method and limits the enhancements for retrieval models when we increase the data size. The data quality from ChatGLM3 is better, but not as good as docT5query. This is because that docT5query has been finetuned on this dataset while ChatGLM3 is used in a zero-shot manner. Moreover, we use a 6B ChatGLM3 instead of a very large model, which may also result in its sub-optimal performance. Among all these methods, human annotations lead to the best-performing models. Therefore, there is still a large room for improvment about using large language models to generate pseudo training data.
4.4. Model-Data Joint Laws
We combine the above observations into a single function that can characterize the joint effects of model size and data size. Inspired by the scaling laws of LLMs (Kaplan et al., 2020), we employ the following equation to describe the scaling effect:
where represents the model size and data size, respectively, and are coefficients. We employ results with different model sizes and data sizes to estimate the coefficients. Figure 6 illustrates the actual contrastive entropy and the predictions. In this figure, solid dots represent the data used for curve fitting, while the dashed line indicates the resulting fitted curve. The red stars denote data points utilized to evaluate the accuracy of our predictions. We can see that the predictions relatively are close to the real values.
@@ 5. Application in Budget Allocation
In this section, we showcase a potential application of the scaling laws for dense retrieval observed in our experiments. We use Eq. 8 in this section.
We attempt to estimate the comprehensive cost associated with the lifecycle of dense retrieval models, including data annotation, model training, and model inference. The total cost of training a model with parameters using data points is given by:
@@
Now we estimate the approximate values for . The cost of human annotations is approximated at $0.6 per query-passage pair (Althammer et al., 2023). For computational costs, according to previous studies (Kaplan et al., 2020; Clark et al., 2020), the training and inference computation for Transformer can be assumed by and FLOPs, respectively. We refer to common cloud computing and the price for using an A100 80G GPU is assumed to be $3.93 per hour333https://cloud.google.com/compute/gpus-pricing, with the peak computational power around 312 TFLOPs. For the training phase, we assume that the model is trained for 10,000 steps on a single A100 GPU. At each step, the model encodes a query, a positive passage and a negative passage with a batch size of 256. Each query is around 30 tokens and each passage is around 60 tokens. For the inference phase, we assume that the model is employed in a web search engine. Based on public statistics, we assume that there are around 30 trillion web pages in Google’s index444From https://en.wikipedia.org/wiki/Google_Search, the estimated size of Google’s index is around 30 trillion in 2012.. The inference cost for a dense retrieval model predominantly involves encoding the entire corpus. We estimate that each web page contains approximately 512 tokens. We assume the GPU utilization efficiency is 25%, then we have
@@
We first excludes the cost of inference and only focuses on annotation and training. Figure 7 shows the predicted contrastive entropy against model size under different cost budget. It is clear that for a fixed cost budget, as the training model size increases, the predicted retrieval performance initially improves and then slowly degenerates. The improvement is because that a relatively larger model is more data-efficient and can exhibit stronger ranking performance than smaller models. Nevertheless, when models are too large, only limited budget can be used for data annotation. The data limitations make the performance degenerate. Overall, if we do not consider inference cost, for a 20,000$ budget, it is optimal to train a model with 13 billion parameters. This is primarily due to that larger models are more data-efficient and that human annotation is significantly more expensive than training. Therefore, under a limited budget, maximizing model size can yield better results.
We further include the inference costs into this analysis. The result is shown in Figure 8. It is clear that the optimal model size significantly decreases to only million-scale parameters, even under a larger budget. This is because that the inference cost is huge compared to training cost () and that the small models are more inference-efficient. A billion-scale model will make the inference cost prohibitively high.
6. Limitation and Future Work
This study pioneers the investigation of scaling laws in dense retrieval. We cover major factors like model scale, datasets, training volume, and annotation methods. Several other aspects remain unexplored and can be addressed in future research.
Our experiments utilize contrastive entropy as the evaluation metric due to its continuity, which addresses the discrete nature of ranking metrics and facilitates the derivation of scaling laws. Although we demonstrate a positive correlation between contrastive entropy and ranking performance, it is important to note that they are not equivalent. For instance, as shown in Figure 2, similar contrastive entropy scores do not guarantee similar ranking performance. Future research may explore alternative metrics that might offer a more direct correlation with ranking outcomes.
The training process in this work is based on random negative sampling and contrastive learning. We do not cover more sophisticated training techniques, such as hard negative sampling (Xiong et al., 2021; Zhan et al., 2021), distillation (Hofstätter et al., 2021), and contrastive pre-training (Izacard et al., 2022; Gao and Callan, 2022, 2021). These methods could potentially influence the scaling behaviors observed and should be investigated in the future.
We focus on a common dense retrieval architecture where text is mapped to a single dense vector of fixed dimensionality. However, some researchers have experimented with variations in this architecture, such as mapping to vectors of varying dimensions (Reimers and Gurevych, 2021), multiple vectors (Luan et al., 2021; Khattab and Zaharia, 2020), or even sparse vectors (Formal et al., 2021b; Mallia et al., 2021). Future work could explore how these architectural modifications impact the scaling laws for dense retrieval.
Our evaluations are conducted within in-domain datasets. Although we also attempt out-of-domain test (not reported in the paper), the available datasets are relatively small and yield unstable results, making it challenging to draw robust conclusions. Thus, our results do not currently account for out-of-domain scenarios, and more extensive evaluation could be beneficial in future work.
While we try to assess scaling across various scales, our resources limit the maximum size of our models and the extent of data size. Future work could further evaluate scaling laws on an even larger scale with more extensive models and annotations.
7. Conclusion
This paper systematically investigates the scaling laws of dense retrieval. We conduct experiments on both Chinese and English datasets to assess the impact of model size, data size, and annotation methods on retrieval performance. By utilizing contrastive entropy as the metric, we observe a power law relationship between performance and both model size and data size across different annotation methods and datasets. We also show that the scaling laws help optimize training processes. For instance, the scaling laws is important to budget allocation management, as demonstrated in our experiments. Moreover, scaling laws allows to evaluate the efficacy of different annotation methods. As shown in our experiments, there is still a large improvement room for using large language models to generate relevance annotations. We believe scaling laws offer a systematic approach to assess and improve the training processes of ranking models. While this study has laid a foundation for future exploration in this area, further research is needed to expand our understanding of scaling laws across more varied domains, scales, architectures, and evaluations.
Appendix A Appendix
We use the following prompt for ChatGLM3 to generate queries for one passage. Note that {} is the placeholder for the actual passage.